# Entropy-regularized optimal transport over networks with incomplete marginals information [^1] Supported by NYUAD grant 76/71260/ADHPG/AD356/51119

### v1

@@ I Introduction
Problems of optimal transport over networks have been studied in the past fifteen years from several different angles, see e.g pavon2010discrete ; BCCNP ; chen2016robust ; leoD ; chen2017efficient ; Pey19 ; Z ; HCK ; CLZ ; DB ; HRCK ; KDO ; LCFS ; WFBR ; CGKS . Consider a road or computer network modeled as a directed graph where the nodes represent cities or junctions. Suppose the total number of traveling vehicles/packets is known during a discrete time interval $[0,N]$. Suppose also that at $t=0$ the vehicle/packets distribution can only be observed at certain nodes and similarly at the final time $N$. We are interested in determining the flow on $[0,N]$. Given that the problem is clearly ill-posed, we ask: Can we at least determine what flow is most likely? In spite of the apparent importance, this problem has so far been ignored in the literature. Indeed, on the subsets where the marginal distribution is known, it is natural to impose a hard constraint. This is different from adding divergences to the cost function as done in Problem 3.1 in liero2018optimal , and in chizat2018scaling .
@@ III.1 Half-bridges with partial information
Let $\mathcal{P}(\rho_{N})$ be the family of distributions in $\mathcal{P}$ whose initial marginal $q_{N}$ coincides with $\rho_{N}$ on $\mathcal{X}_{N}$, i.e. $\left[q_{N}(x)-\rho_{N}(x)\right]\mathds{1}_{\mathcal{X}_{N}}\equiv 0$. By the same argument as with full final marginal knowledge, the minimizer of ${{\mathbb{D}}}(Q\|P)$ over $\mathcal{P}(\rho_{N})$ is the distribution with $\bar{q}^{*}(N,x_{N};0,x_{0})=\bar{p}(N,x_{N};0,x_{0})$ and minimizing ${{\mathbb{D}}}(q_{N}\|p_{N})$. By the same argument as before, the problem reduces to minimize $\sum_{x_{N}\in\mathcal{(}X_{N})^{c}}\log\frac{q_{N}(x_{N})}{p_{N}(x_{N})}q_{N}(x_{N})$ over $q_{N}$ coinciding with $\rho_{N}$ on $\mathcal{X}_{N}$ subject to the constraint
@@ III.3 An iterative algorithm
We thus find a value for $c_{0}$ namely
@@ IV.2 Ruelle-Bowen prior
when the prior is given by the Ruelle-Bowen (RB) random walk measure Rue04 ; DL . This measure induces a uniform probability measure on paths of a fixed length between any two given nodes. Thus, the Ruelle-Bowen measure represents a natural choice for the prior in cases where we only know the topology of the graph. This distribution maximises the entropy rate for a random walker. The latter is bounded by the topological entropy rate which depends only on the topology of the graph and not its underlying probability distribution.
@@ VII Conclusion
We have shown that regularized optimal transport over networks can be solved also in the case of partial information about the initial and final distribution. Two cases were considered: In case one, one knows one or both of the marginals only on a subset of the state space, see Section III. In the second case, only some moments of the initial and/or final marginal are known, see Section VI.The solution in the first case can be computed through an efficient iterative scheme which represents a suitable modification of the classical Fortet-IPF-Sinkhorn algorithm. In the second case, standard convex optimization iterative methods can be used to solve the dual problem. Similar results in both cases can be established in the continuous-time continuous state space case, see Appendix C. Albeit the mathematics is admittedly simple, the potential of this result appears considerable given that the solution also entails a most likely completion of the initial and final marginals.

### v2

@@ I Introduction
Problems of optimal transport over networks have been studied in the past fifteen years from several different angles, see e.g pavon2010discrete ; BCCNP ; chen2016robust ; leoD ; chen2017efficient ; Pey19 ; Z ; HCK ; CLZ ; DB ; HRCK ; KDO ; LCFS ; WFBR ; CGKS . Consider a road or computer network modeled as a directed graph where the nodes represent cities or junctions. Suppose the total number of traveling vehicles/packets is known during a discrete time interval $[0,N]$. Suppose also that at $t=0$ the vehicle/packets distribution can only be observed at certain nodes and similarly at the final time $N$. We are interested in determining the flow on $[0,N]$. Given that the problem is clearly ill-posed, we ask: Can we at least determine what flow is most likely? In spite of the apparent importance, this problem has so far received limited attention in the literature. In the recent paper Mascherpa2023 , authors develop an application to estimating the pollution spreading in a network as a discrete multi-marginal Schrödinger bridge where the information on the intermediate marginals is only on a fixed subset of the state space. Full observation of the initial marginal is however assumed. The dual problem is solved in Mascherpa2023 through a block coordinate ascent method. While we study the simpler problem with initial-final marginals, we do not assume full knowledge of the initial marginal. Thus, our solution entails also the most likely completion of the initial and final marginals. Moreover, the two marginals may be known on different subsets of the state space. The role of the initial marginal of the prior distribution on paths, which is missing in Mascherpa2023 , is clarified in this paper. Moreover, the normalization condition of the marginal has to be explicitly imposed in our formulation of the problem. Without that, the solution does not admit an immediate large deviation (most likely) interpretation (see, however, e.g. chen2021most ). Indeed, in the simulations of Section IV, without the normalization condition the mass of traveling on the network does not sum to one. Notice that, on the subsets where the marginal distribution is known, it is natural to impose a hard constraint. This is different from adding divergences to the cost function as done in Problem 3.1 in liero2018optimal , and in chizat2018scaling .
@@ III.1 Half-bridges with partial information
Let $\mathcal{P}(\rho_{N})$ be the family of distributions in $\mathcal{P}$ whose final marginal $q_{N}$ coincides with $\rho_{N}$ on $\mathcal{X}_{N}$, i.e. $\left[q_{N}(x)-\rho_{N}(x)\right]\mathds{1}_{\mathcal{X}_{N}}\equiv 0$. By the same argument as with full final marginal knowledge, the minimizer of ${{\mathbb{D}}}(Q\|P)$ over $\mathcal{P}(\rho_{N})$ is the distribution with $\bar{q}^{*}(N,x_{N};0,x_{0})=\bar{p}(N,x_{N};0,x_{0})$ and minimizing ${{\mathbb{D}}}(q_{N}\|p_{N})$. By the same argument as before, the problem reduces to minimize $\sum_{x_{N}\in\mathcal{(}X_{N})^{c}}\log\frac{q_{N}(x_{N})}{p_{N}(x_{N})}q_{N}(x_{N})$ over $q_{N}$ coinciding with $\rho_{N}$ on $\mathcal{X}_{N}$ subject to the constraint
@@ III.3 An iterative algorithm
We thus find the value for $c_{0}$, namely
@@ IV.2 Ruelle-Bowen prior
when the prior is given by the Ruelle-Bowen (RB) random walk measure Rue04 ; DL , (chen2016robust, , SectionIV). This measure induces a uniform probability measure on paths of a fixed length between any two given nodes. Thus, the Ruelle-Bowen measure represents a natural choice for the prior in cases where we only know the topology of the graph. This distribution maximises the entropy rate for a random walker. The latter is bounded by the topological entropy rate which depends only on the topology of the graph and not its underlying probability distribution.
@@ VII Conclusion
We have shown that regularized optimal transport over networks can be solved also in the case of partial information about the initial and final distribution. Two cases were considered: In case one, one knows one or both of the marginals only on a subset of the state space, see Section III. In the second case, only some moments of the initial and/or final marginal are known, see Section VI.The solution in the first case can be computed through an efficient iterative scheme which represents a suitable modification of the classical Fortet-IPF-Sinkhorn algorithm. In the second case, standard convex optimization iterative methods can be used to solve the dual problem. Similar results in both cases can be established in the continuous-time continuous state space case, see Appendix C. The problem with moment information has been studied in the continuous setting without entropy regularization in Alfonsi2021 . In Elvander2020 , a multi-marginal regularized optimal transport problem is studied where the partial information on the marginals is available through various linear maps. Albeit the mathematics is admittedly simple, the potential of this result appears considerable given that the solution also entails a most likely completion of the initial and final marginals.